Beth is seven years older than Una. She is twice as old as Una was when Beth was as old as Una is now. How old is each now?
I have tried doing B = U+7 and B-7 = 2U, but it doesn't work. How do you represent the second part?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52781)

.

Right equations are


    B = U + 7

    B = 2*(U-7).


When you solve it, you will get the ANSWER  U = 21,  B = 28.

Answered.



Answer by MathTherapy(10552)

Beth is seven years older than Una. She is twice as old as Una was when Beth was as old as Una is now. How old is each now?
I have tried doing B = U+7 and B-7 = 2U, but it doesn't work. How do you represent the second part?
Let Beth’s age, be B
Difference in ages: 7
Beth was Una’s age, 7 years ago, or she was: B - 7, at which time, Una was B - 7 - 7 = B - 14
We then get: B = 2(B - 14)
That's the ONLY equation you need, which you can use to determine B's (Beth's) age.